OpenStudy (anonymous):

What is the solution set for |-2x + 16| - 9 = -3 A. x = -5 and x = -11 B. x = 5 and x = 11 C. x = -5 and x = 11 D. x = 5 and x = -11

OpenStudy (anonymous):

First, isolate the abs value expression on one side, like this |-2x + 16| - 9 = -3 +9 +9 |-2x + 16| = 6 Now, the stuff inside the abs value has to = 6 or = -6 Set up 2 equations and solve each one to get your 2 answers.